# Why Can’t I Open My Drawer? Mitigating Object-Driven Shortcuts in Zero-Shot Compositional Action Recognition

### v1

@@ Abstract
We study Compositional Video Understanding (CVU), where models must recognize verbs and objects and compose them to generalize to unseen combinations. We find that existing Zero-Shot Compositional Action Recognition (ZS-CAR) models fail primarily due to an overlooked failure mode: object-driven verb shortcuts. Through systematic analysis, we show that this behavior arises from two intertwined factors—severe sparsity and skewness of compositional supervision, and the asymmetric learning difficulty between verbs and objects. As training progresses, the existing ZS-CAR model increasingly ignores visual evidence and overfits to co-occurrence statistics. Consequently, the existing model does not gain the benefit of compositional recognition in unseen verb-object compositions. To address this, we propose RCORE, a simple and effective framework that enforces temporally grounded verb learning. RCORE introduces (i) a composition-aware augmentation that diversifies verb–object combinations without corrupting motion cues, and (ii) a temporal order regularization loss that penalizes shortcut behaviors by explicitly modeling temporal structure. Across two benchmarks—Sth-com and our newly constructed EK100-com—RCORE significantly improves unseen composition accuracy, reduces reliance on co-occurrence bias, and achieves consistently positive compositional gaps. Our findings reveal object-driven shortcuts as a critical limiting factor in ZS-CAR and demonstrate that addressing them is essential for robust compositional video understanding. The code is available at https://github.com/KHU-VLL/RCORE.
(a) Co-occurrence bias
(b) Asymmetric learning difficulty
Figure 1: Why object-driven shortcuts emerge in compositional video understanding? (a) Co-occurrence bias. Datasets are intrinsically sparse and highly skewed in their verb–object combinations, creating strong co-occurrence priors. Models exploit these priors as a shortcut: once the object is recognized, the model often predicts the most frequent verb paired with it, ignoring temporal evidence. (b) Asymmetric learning difficulty. Objects are visually explicit and easy to recognize from a single frame, whereas verbs require multi-frame temporal reasoning. This imbalance makes object features dominate training signals. Together, these two factors drive models to adopt object-driven shortcuts, hindering genuine compositional generalization. Best viewed with zoom and color.
11footnotetext: This work was done during an internship at NAVER Cloud.22footnotetext: Corresponding author.
@@ 1 Introduction
Many human actions can be decomposed into two semantic primitives—verbs and objects—and robust video understanding requires recognizing each component and reasoning over their interaction as a composition [12, 22, 7, 8, 2]. Zero-Shot Compositional Action Recognition (ZS-CAR) formalizes this goal: a model must correctly identify unseen verb–object combinations despite only being exposed to a sparse subset of compositions during training [16].
However, existing ZS-CAR models exhibit a characteristic and persistent failure: they mispredict verbs in unseen compositions by relying on object-driven shortcuts. In other words, rather than learning verb semantics from temporal evidence, models exploit objects as the dominant cue whenever such shortcuts correlate with training labels. To the best of our knowledge, this failure mode has not been explicitly identified or systematically addressed in prior ZS-CAR approaches [16, 14]. We show that this shortcut failure mode arises from two intertwined issues of the ZS-CAR setting: (i) Sparsity and skewness of compositional supervision. The combinatorial nature of verb–object pairs yields extreme under-coverage of the compositional space, e.g. only $\sim 14.0\%$ label coverage ratio in the Something-Else [22] and EPIC-KITCHENS-100 [7] datasets. This induces strong co-occurrence statistics to the deep learning model as depicted in Figure 1 (a). (ii) Asymmetric learning difficulty between verbs and objects. Objects are recognizable from single frames, while verbs require multi-frame temporal reasoning, as shown in Figure 1 (b). This imbalance naturally encourages models to treat objects as easy predictive shortcuts [25, 17, 32, 11, 30], especially under sparse data regimes.
Our analysis reveals that the existing ZS-CAR model amplifies these issues during training. They increasingly bias predictions toward frequently co-occurring compositions, exhibit growing false-seen and false-co-occurrence prediction ratios, and ultimately yield negative compositional gaps—indicating that they fail to extract any benefit from compositional reasoning beyond independent verb and object recognition. (See Section 3 for details.) These observations call into question the robustness of current ZS-CAR approaches and motivate the need for techniques that explicitly target shortcut mitigation.
We propose RCORE (Robust COmpositional REpresentations), a simple and effective framework designed to mitigate object-driven shortcut and promote temporally grounded verb representations. RCORE consists of two components: (i) a composition-aware augmentation that synthesizes plausible unseen verb–object combinations while preserving the temporal structure of the original video; and (ii) a temporal order regularization objective that penalizes reliance on static cues by contrasting original and perturbed temporal representations. Together, these components address the two root causes of failure in ZS-CAR: compositional sparsity and asymmetric learning difficulty.
Across both datasets, RCORE significantly strengthens verb recognition, reduces co-occurrence bias, produces positive compositional gaps, and improves unseen composition accuracy over strong conditional baselines. Our findings indicate that the primary obstacle in compositional video understanding is not model capacity but model behavior: ZS-CAR models fail because they resort to object-driven shortcuts. By directly tackling this failure mode, RCORE provides a simple path toward more reliable and generalizable ZS-CAR.
Figure 2: Controlled experiments demonstrate object-driven shortcut learning in ZS-CAR. We empirically identify a key failure mode in ZS-CAR—object-driven shortcuts. (a) Objects are easier to learn than verbs. We train a randomly initialized ViT [10] on a balanced 10 $\times$ 10 verb-object subset from Sth-com [16]. The learning curves show that object accuracy increases much faster than verb accuracy, indicating that objects are easier to learn. (b) Under a strong co-occurrence bias, models exploit object cues as shortcuts, failing to learn robust verb representations. We prepare a perfectly biased training set and two kinds of test splits utilizing the Sth-com [16] dataset: seen compositions (bias-aligned) and unseen compositions (bias-conflict). Both the randomly initialized ViT [10] and pretrained CLIP [28] exhibit a consistent trend: while unseen object accuracy increases alongside seen performance, unseen verb accuracy concurrently drops below random chance. Best viewed with zoom and color.
@@
  • Identifying the core failure mode of ZS-CAR. We show that existing models rely on object-driven verb shortcuts, caused by the sparsity of compositional supervision and the asymmetric learning difficulty between verbs and objects—an issue not explicitly recognized in prior work.
  • Diagnosing limitations of the ZS-CAR evaluation protocol. We reveal that common practices, such as test-set bias calibration, obscure true compositional generalization. To address this, we adopt an open-world unbiased evaluation setting and introduce the compositional gap metric to measure the benefit of compositional reasoning.
  • A simple yet effective framework that mitigates shortcut learning. We propose RCORE, which comprises a composition-aware augmentation and a temporal order regularization loss (TORC) to promote temporally grounded verb representations under compositional sparsity. Through experiments on Sth-com and our new EK100-com dataset, we demonstrate that RCORE shows consistently favorable performance improvements across diverse evaluation scenarios, metrics, and datasets.
@@ 2 Related Work
Zero-shot compositional action recognition. Zero-Shot Compositional Action Recognition (ZS-CAR) [16] aims to generalize to novel verb–object combinations not seen during training. Unlike zero-shot (or open-set) action recognition [5, 4, 40], ZS-CAR retains the same verb and object classes across train and test splits and evaluates a model’s ability to compose known primitives. It addresses the performance degradation issue on novel unseen combinations of known concepts, which arises from the model’s lack of compositional generalization ability. Conditional learning [16, 13] has been effective for estimating joint verb–object probabilities, while other work [14] pursues improved feature disentanglement. However, as our analysis reveals in Section 3 and Section 5.2, these approaches remain susceptible to object-driven shortcut behavior and progressively overfit co-occurrence statistics.
Compositional zero-shot Learning in image understanding.
Compositional Zero-Shot Learning (CZSL) has been widely studied in the image domain [23, 24, 21, 34, 36, 26]. CZSL enables models to recognize novel combinations of previously learned primitive concepts, such as objects and their states. Prior works have approached this problem through feature disentanglement [29, 42], modular architectures [27], and CLIP-based prompt learning [26, 18]. Unlike image-based CZSL, ZS-CAR requires modeling temporal dynamics for verbs, making the problem fundamentally different. Our work focuses specifically on the video domain and highlights a failure mode—object-driven verb shortcuts—that does not arise in image-only CZSL.
Shortcut learning.
Shortcut learning is a pervasive issue in visual recognition. Several studies show that models tend to rely on easy-example biases [35, 15, 30] or easier-to-learn cues—such as color or texture—when these cues correlate with labels [25, 11]. A line of work attributes this behavior to co-occurrence statistics in the training data [1, 32]. In video understanding, similar shortcut behaviors occur when action models rely on backgrounds rather than motion [17, 6, 33, 9, 2]. Building on these observations, we argue that ZS-CAR introduces an additional structural imbalance: verbs require temporal reasoning, whereas objects can be recognized from a single frame. This learning asymmetry amplifies shortcut tendencies and leads models to predict verbs primarily from object cues—a failure mode we diagnose and directly address in this work.
@@ 3.1 Zero-Shot Compositional Action Recognition
We study Zero-Shot Compositional Action Recognition (ZS-CAR) [16], where the goal is to recognize verb–object compositions that might never appear in the training set. Let $\mathbb{Y}^{V}$ and $\mathbb{Y}^{O}$ denote the sets of verb and object labels, and let each action label be a pair $\mathbf{y}=(\mathbf{y}^{V},\mathbf{y}^{O})\in\mathbb{Y}^{V}\times\mathbb{Y}^{O}$. The training set is $\mathbb{D}_{\text{train}}=\{(\mathbf{X}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$, and its set of observed (seen) compositions is $\mathbb{Y}_{\text{seen}}=\{\mathbf{y}_{i}\mid(\mathbf{X}_{i},\mathbf{y}_{i})\in\mathbb{D}_{\text{train}}\}$. The space of unseen compositions is $\mathbb{Y}_{\text{unseen}}=(\mathbb{Y}^{V}\times\mathbb{Y}^{O})\setminus\mathbb{Y}_{\text{seen}}$. We define the validation and test sets analogously. The validation and test sets contain both seen and unseen compositions, enabling evaluation of compositional generalization. ZS-CAR therefore measures whether a model can correctly compose familiar verbs and objects into novel pairs at test time.
@@ 3.2 Diagnostic Metrics
For better understanding the underlying behavior of ZS-CAR models, we present two diagnostic metrics: (1) training bias metrics to measure a model’s overfitting and (2) Compositional Gap to assess the benefits of compositional modeling.
Training bias metrics
To measure how the model overfits to the training set statistics, we introduce two types of misclassification ratios: False Seen Prediction and False Co-occurrence Prediction ratios. False Seen Prediction (FSP) ratio denotes the ratio of unseen composition samples incorrectly predicted to seen composition labels. Out of these mispredicted samples, False Co-occurrence Prediction (FCP) ratio is likewise defined as the proportion of those assigned to frequently co-occurring labels. Here, a composition label is considered ‘frequently co-occurring’ if its occurrence is higher than one standard deviation above the mean frequency in the training data.
Compositional Gap
We introduce Compositional Gap to examine whether a model genuinely benefits from understanding videos in a compositional manner. Compositional Gap ($\Delta_{\text{CG}}$) is defined as
$\Delta_{\text{CG}}=\mathrm{Acc}^{C}-(\mathrm{Acc}^{V}\times\mathrm{Acc}^{O}),$ (1)
where $\text{Acc}^{C}$ is the composition accuracy (%) computed from the logits over the composition space, whereas $\text{Acc}^{V}$ and $\text{Acc}^{O}$ are the verb and object accuracies (%) derived from the logits over their individual component spaces. A positive $\Delta_{\text{CG}}$ indicates that the model truly gains an advantage from its compositional architecture.
Figure 3: Learning curve of the SOTA model with our diagnostic metrics. We plot the learning curve of C2C [16] trained on Sth-com [16]. We measure the False Seen Prediction (FSP) and False Co-occurrence Prediction (FCP) ratios, and observe that the seen–unseen accuracy gap ($\Delta_{SU}$) correlates strongly with both metrics. There observations indicate that the current SOTA model exhibits overfitting to seen compositions. Best viewed with zoom and color.
Figure 4: Overview of RCORE. (a) Overview of our proposed RCORE framework. (b) VOCAMix synthesizes plausible yet unseen verb–object compositions while preserving the temporal structure of the primary video. (c) TORC penalizes alignment between original and temporally perturbed feature vectors, enforcing explicit temporal order modeling and reducing object-driven shortcuts.

### v2

@@ Abstract
Zero-Shot Compositional Action Recognition (ZS-CAR) requires recognizing novel verb–object combinations composed of previously observed primitives. In this work, we tackle a key failure mode: models predict verbs via object-driven shortcuts (i.e., relying on the labeled object class) rather than temporal evidence. We argue that sparse compositional supervision and verb–object learning asymmetry can promote object-driven shortcut learning. Our analysis with proposed diagnostic metrics shows that existing methods overfit to training co-occurrence patterns and underuse temporal verb cues, resulting in weak generalization to unseen compositions. To address object-driven shortcuts, we propose Robust COmpositional REpresentations (RCORE) with two components. Co-occurrence Prior Regularization (CPR) adds explicit supervision for unseen compositions and regularizes the model against frequent co-occurrence priors by treating them as hard negatives. Temporal Order Regularization for Composition (TORC) enforces temporal-order sensitivity to learn temporally grounded verb representations. Across Sth-com and EK100-com, RCORE reduces shortcut diagnostics and consequently improves compositional generalization.
This work was done during an internship at NAVER Cloud.Corresponding author.
@@ 1 Introduction
Many human actions can be decomposed into two semantic primitives—verbs and objects—and robust video understanding requires recognizing each component and reasoning over their interaction as a composition [goyal2017something, mat2020sthelse, damen2022ek100, diba2020hvu, bae2024devias]. Zero-Shot Compositional Action Recognition (ZS-CAR) formalizes this goal: a model must recognize unseen verb–object pairs while keeping the verb and object vocabularies shared across splits [li2024c2c, jiang2025dhd, jung2025crr, ye2025logic].
In this setting, a key failure mode emerges: models often predict verbs via object-driven shortcuts rather than temporal evidence. Throughout, object refers to the labeled noun class in the verb–object vocabulary (not unlabeled background entities). We argue that object-driven shortcuts arise from two issues inherent to ZS-CAR.
Figure 1: Why object-driven shortcuts emerge in compositional video understanding? (a) Co-occurrence bias. Datasets are intrinsically sparse and highly skewed in their verb–object combinations, creating strong co-occurrence priors. (b) Asymmetric learning difficulty. Objects are visually explicit and easy to recognize from a single frame, whereas verbs require multi-frame temporal reasoning. This imbalance makes object features dominate training signals. (c) Object-driven shortcuts. Together, these two factors drive models to adopt object-driven shortcuts, hindering genuine compositional generalization. Once the object is recognized, the model often predicts the most frequent verb paired with it, ignoring temporal evidence.
First, compositional supervision is sparse and skewed. As illustrated in Figure 1 (a), datasets cover only a small fraction of the combinatorial verb–object space, and training labels concentrate on a limited set of frequent pairs [mat2020sthelse, damen2022ek100]. This skew induces a strong co-occurrence prior in the training data—the empirical dominance of frequent verb–object pairs—which can be amplified into a model shortcut at training time. Under such skewed supervision, predictions for unseen compositions tend to drift toward seen (especially frequent) pairs. Second, verbs are harder to learn than objects. As shown in Figure 1 (b), an object often becomes recognizable from a single frame, whereas a verb typically requires multi-frame temporal reasoning. Prior work on shortcut learning suggests that models tend to rely on easier-to-learn cues rather than the intended evidence [nam2020lff, li2018resound, singh2020dont, geirhos2018cuecon, scimeca2022shortcut]. In ZS-CAR, this verb–object difficulty gap makes object cues an especially attractive shortcut for verb prediction.
In this work, we posit that these two intertwined issues in verb-object compositional learning promote object-driven shortcuts, which we diagnose as the key factor hindering ZS-CAR. As Figure 1 (c) illustrates, a model can match a verb once it recognizes its frequently co-occurring object during training, without modeling the temporal dynamics essential to correctly recognize the verb. This behavior weakens generalization to unseen compositions because it ties verb prediction to object-dependent priors rather than temporally grounded evidence. We observe that models [li2024c2c], even with a modern video-pretrained backbone [wang2024iv2], indeed confuse verbs with opposite temporal order (e.g., opening vs. closing) on unseen compositions (see Section 5.2).
To dissect how object-driven shortcuts hinder unseen generalization, we perform a comprehensive quantitative diagnosis. First, through controlled experiments, we show that the asymmetric learning difficulty between verbs and objects, together with skewed co-occurrence priors, promotes verb shortcut learning driven by object cues. Next, we quantify shortcut-driven failures in existing ZS-CAR models by two diagnostic ratios: False Seen Prediction (FSP) and False Co-occurrence Prediction (FCP), which measure how often an unseen input collapses to a seen—and especially frequent—training composition. Using FSP/FCP, we show that a strong ZS-CAR baseline [li2024c2c] remains heavily biased toward seen compositions, resulting in poor generalization to unseen compositions.
Driven by the diagnosis, we propose Robust COmpositional REpresentations (RCORE), targeting the two root causes of object-driven shortcuts. First, Co-occurrence Prior Regularization (CPR) expands supervision over originally absent compositions and suppresses the dominance of frequently seen pairs by treating them as hard negatives. Second, Temporal Order Regularization for Composition (TORC) enforces temporal-order sensitivity so that the model learns temporally grounded verb representations instead of relying on static cues.
We validate RCORE on Sth-com [li2024c2c] under a realistic open-world evaluation protocol that evaluates all possible verb-object pairs, without requiring test-set ground truth labels. We also introduce EK100-com, a ZS-CAR dataset repurposed from EPIC-KITCHENS-100 [damen2022ek100], which exhibits more severe compositional sparsity than Sth-com [li2024c2c]. Across datasets and backbones, RCORE reduces shortcut diagnostics (FSP/FCP) and improves unseen compositional generalization.
@@
  • Shortcut Diagnosis: We provide a diagnostic view of object-driven shortcuts in ZS-CAR. We empirically identify strong co-occurrence priors and asymmetric learning difficulty as the root causes of the shortcuts. Furthermore, we expose the limitations of baselines by quantifying how unseen inputs collapse into seen-especially frequent-training pairs, using FSP/FCP.
  • Diagnosis-Driven Framework: We propose RCORE comprising CPR to suppress co-occurrence priors by expanding supervision over originally absent compositions and TORC to promote temporally grounded verb learning.
  • Effectiveness & Robustness: We demonstrate that RCORE is effective, achieving superior performance on unseen compositions over baselines on Sth-com and EK100-com datasets, with multiple VLM backbones. Moreover, through extensive quantitative analyses, we prove that our approach mitigates co-occurrence biases and learns temporally grounded verb features.
@@ 2 Related Work
Compositional Action Recognition (CAR). CAR aims to recognize actions by factoring them into verbs and objects. A recurring challenge is learning verb representations that are robust to static cues (e.g. objects/scenes) and capture temporal dynamics [mat2020sthelse, sun2021cdn, chatterjee2023oap]. Prior work addresses this by modeling object-level interactions [mat2020sthelse, sun2021cdn] or by training verb representations to be less sensitive to object appearance [chatterjee2023oap]. Related directions include open-vocabulary CAR, which expands the action label space beyond a fixed taxonomy (often via text-driven semantics) [luo2022dark, chatterjee2023oap]. In this work, we study Zero-Shot Compositional Action Recognition (ZS-CAR) under a fixed verb/object vocabulary, where verb and object label sets are shared across splits but some verb–object pairs are held out for testing [li2024c2c, jiang2025dhd, jung2025crr, ye2025logic]. In ZS-CAR, conditional learning [li2024c2c, jiang2025dhd] and disentanglement-oriented designs [jung2025crr] improve joint modeling, yet our study highlights that these models can still exhibit object-driven shortcut behavior under sparse and skewed compositional supervision (see Section 3, Section 5.2).
Compositional Zero-Shot Learning (CZSL) for image understanding. In the image domain, CZSL focuses on recognizing novel compositions of seen primitives (e.g., attribute–object) under sparse compositional supervision [misra2017red, naeem2021czsl, mancini2021compcos, kim2023hierarchical, wu2025cond, nayak2023csp]. Representative approaches include feature disentanglement [saini2022disen, zhang2024disen], modular factorization [purushwalkam2019taskdriven], and CLIP-based prompt learning [nayak2023csp, li2024context]. Compared to image CZSL, ZS-CAR additionally requires temporal grounding of verbs, which changes both the failure modes and the regularization opportunities. In contrast, we reveal how sparse compositional supervision, coupled with the higher learning difficulty of verbs, amplifies shortcut reliance in videos, and propose training-time regularizers tailored to ZS-CAR.
Shortcut learning. Models can learn shortcuts by exploiting spurious correlations that are easier to fit than the intended evidence [geirhos2018cuecon, ilyas2019adversarial, nam2020lff, scimeca2022shortcut]. Such shortcuts are closely tied to skewed training distributions and co-occurrence statistics, which can induce reliance on dataset priors [agrawal2018dont, singh2020dont]. In compositional settings, prior work reweights rare combinations to mitigate co-occurrence bias [kim2023hierarchical]. In video understanding, several works report reliance on static cues such as scenes [whycantchoi, wang2021be, ding2022fame, bae2024devias] and objects [bahng2019rebias, chatterjee2023oap] instead of motion-sensitive evidence. Building on these observations, we study a ZS-CAR-specific manifestation: object-driven shortcuts where object cues dominate verb prediction under sparse verb–object supervision, and we quantify this behavior with dedicated diagnostics.
Vision-Language Model (VLM). Recent VLMs broadly follow two paradigms: encoder-based and decoder-based. CLIP [radford2021clip] is a representative encoder-only VLM pretrained via large-scale contrastive learning, and it exhibits strong zero-shot recognition [radford2021clip] as well as non-trivial compositional behavior [abbasi2024decipher]. Following CLIP, recent works pretrain encoder-based video VLMs on large-scale video-text data [ma2022xclip, li2023unmasked, zhao2024videoprism, wang2024iv2]. Decoder-based VLMs (e.g., LLaVA [liu2023llava], Qwen-VL [bai2025qwen25vl]) couple a vision encoder with an LLM decoder and excel at open-ended generation. In ZS-CAR, however, inference and evaluation operate over a fixed verb–object label space with explicit compositional scores (i.e., logits over $\mathbb{Y}^{V}\times\mathbb{Y}^{O}$), which provides a controlled interface for diagnosis and for designing training-time regularizers that are transferable across backbones. Such decoder-based VLM interfaces often do not expose, in a standardized and reproducible way, explicit scores over the fixed label space, as they primarily provide language outputs at inference time. Therefore, we focus on encoder-based VLMs in this controlled setting, and show that object-driven shortcuts persist even with a video-pretrained backbone [wang2024iv2] when paired with a strong ZS-CAR baseline [li2024c2c]; we then propose RCORE to regularize compositional scoring to mitigate these shortcuts.
@@ 3.1 Zero-Shot Compositional Action Recognition
We study Zero-Shot Compositional Action Recognition (ZS-CAR) [li2024c2c], where the goal is to recognize unseen verb–object compositions that are held out from the training set, while the verb and object vocabularies are fixed and shared across train/validation/test splits. Thus, unlike open-set or open-vocabulary settings, ZS-CAR does not introduce unseen verbs or objects at test time—only unseen compositions (verb–object pairs). Let $\mathbb{Y}^{V}$ and $\mathbb{Y}^{O}$ denote the sets of verb and object labels, and let each composition label be a pair $\mathbf{y}^{C}=(\mathbf{y}^{V},\mathbf{y}^{O})\in\mathbb{Y}^{V}\times\mathbb{Y}^{O}$. The training set is $\mathbb{D}_{\text{train}}=\{(\mathbf{X}_{i},\mathbf{y}^{C}_{i})\}_{i=1}^{N}$, and its set of observed (seen) compositions is $\mathbb{Y}_{\text{seen}}=\{\mathbf{y}^{C}_{i}\mid(\mathbf{X}_{i},\mathbf{y}^{C}_{i})\in\mathbb{D}_{\text{train}}\}$. The space of unseen compositions is $\mathbb{Y}_{\text{unseen}}=(\mathbb{Y}^{V}\times\mathbb{Y}^{O})\setminus\mathbb{Y}_{\text{seen}}$. The validation and test sets contain both seen and unseen compositions, enabling evaluation of compositional generalization.
Figure 2: Controlled experiments demonstrate object-driven shortcut. (a) Objects are easier than verbs. A randomly initialized ViT trained on a balanced $10{\times}10$ Sth-com subset learns objects substantially faster than verbs. (b) Co-occurrence bias induces object-driven shortcuts. On a perfectly biased training split, CLIP (with AIM [yang2023aim]) achieves high object accuracy, but verb accuracy on bias-conflict unseen compositions drops below chance, indicating shortcut-driven verb failures.
@@ 3.2 Diagnostic Metrics
To better understand ZS-CAR model behavior, we introduce two diagnostics: (1) training-bias metrics that quantify over-reliance on training co-occurrence statistics, and (2) the Compositional Gap that measures the benefit of modeling verb–object dependence beyond independent recognition.
Training-bias metrics. To quantify reliance on co-occurrence priors, we define two misclassification ratios on the unseen subset of an evaluation split $\mathbb{Y}_{\text{unseen}}$. False Seen Prediction (FSP) is the fraction of misclassified unseen composition samples whose predictions $\hat{\mathbf{y}}^{C}$ fall into the seen composition categories $\mathbb{Y}_{\text{seen}}$. False Co-occurrence Prediction (FCP) is the fraction of misclassified unseen composition samples whose prediction $\hat{\mathbf{y}}^{C}$ falls into the frequent compositions $\mathbb{Y}_{\text{freq}}$. Let $f(\mathbf{y}^{C})$ denote the training frequency of a seen composition $\mathbf{y}^{C}\in\mathbb{Y}_{\text{seen}}$. We define the set of frequent seen compositions $\mathbb{Y}_{\text{freq}}$ as
$\vskip-3.00003pt\mathbb{Y}_{\text{freq}}=\{\mathbf{y}^{C}\in\mathbb{Y}_{\text{seen}}\mid f(\mathbf{y}^{C})>\mu_{f}+\sigma_{f}\},\vskip-3.00003pt$ (1)
where $\mu_{f}$ and $\sigma_{f}$ are the mean and standard deviation of $\{f(\mathbf{y}^{C})\}_{\mathbf{y}^{C}\in\mathbb{Y}_{\text{seen}}}$. To further analyze shortcut patterns, we decompose FSP/FCP into three component-wise cases based on prediction errors: (i) Verb-collapse, where the predicted object is correct but the verb is incorrect; (ii) Object-collapse, where the predicted verb is correct but the object is incorrect; and (iii) Dual-collapse, where both verb and object are incorrect. Using the metrics, we can identify whether co-occurrence bias primarily harms verb or object learning.
Compositional Gap. We introduce Compositional Gap ($\Delta_{\text{CG}}$) to examine whether a model benefits from predicting the joint verb–object composition beyond getting the two components correct independently. Given joint logits over $\mathbb{Y}^{V}\times\mathbb{Y}^{O}$, we obtain the predicted composition $\hat{\mathbf{y}}^{C}=(\hat{\mathbf{y}}^{V},\hat{\mathbf{y}}^{O})$ by selecting the top-1 composition label. We then compute top-1 accuracies $\mathrm{Acc}^{V}$, $\mathrm{Acc}^{O}$, and $\mathrm{Acc}^{C}$ respectively. Then we define $\Delta_{\text{CG}}$ as
$\vskip-3.00003pt\Delta_{\text{CG}}=\mathrm{Acc}^{C}-(\mathrm{Acc}^{V}\times\mathrm{Acc}^{O}).\vskip-3.00003pt$ (2)
We report $\Delta_{\text{CG}}$ alongside FSP/FCP and unseen composition accuracy as a diagnostic metric, rather than interpreting it as a standalone evaluation criterion.
(a) CLIP
(b) InternVideo2
(c) FSP/FCP breakdown
Figure 3: Shortcut reliance correlates with poor unseen generalization. We show learning curves of C2C [li2024c2c] on Sth-com with CLIP [radford2021clip] and InternVideo2 [wang2024iv2]. (a) With CLIP, the seen–unseen accuracy gap grows together with FSP/FCP, indicating strong overfitting to seen compositions. (b) Even with the video-pretrained InternVideo2 backbone, the model still shows a notable seen–unseen gap and high FSP. (c) FSP/FCP breakdown at the final epoch: CLIP is dominated by verb-collapse due to object-driven shortcuts, while InternVideo2 exhibits milder but persistent shortcut behavior. Best viewed with zoom and color.